OpenStudy (anonymous):

Which of the following is true about the relationship between the slopes of the lines whose equations are 8x - 9y = 5 and 9x - 8y = 1?

OpenStudy (anonymous):

They are equal. They are reciprocals. They are opposites.

OpenStudy (anonymous):

First change them into Slope-Intercept form. y = mx + b. m is slope b is y-intercept Then you can compare the slopes! (:

OpenStudy (anonymous):

8x - 9y = 5 and 9x - 8y = 1? for 8x - 9y = 5 8x-5=9y separate y to the other side y=(8/9 )x -(5/9......... EQ 1) and for 9x - 8y = 1 separate y to the other side 9x-1=8y y=(9/8)x - 1/8 ....EQ 2) can you see the difference of the slope on each EQ?

OpenStudy (anonymous):

the difference on the slope of EQ 1) and EQ 2)

OpenStudy (anonymous):

umm no